A die is thrown twice. What is the probability of getting at least one six?A1/3B5/18C11/36D25/36 Correct answer: C ExplanationP(no six in both throws) = (5/6)² = 25/36, so P(at least one six) = 1 − 25/36 = 11/36. Join the discussion Report an error